OpenStudy (anonymous):

i will fan and medal!!!!!!!! The regular price of an item is $65. The item is on sale with a discount of $8. What is the discount rate? A. 0.123% B. 1.23% C. 8% D. 12.3%

563blackghost (563blackghost):

To find the discount rate you would divide the discount over the regular price of the item in which you would then multiply by 100 to find the percent.... \(\Huge{\frac{8}{65}= ? \times 100=discount~rate}\)

OpenStudy (anonymous):

i think the answer is D.

563blackghost (563blackghost):

That would be correct ^^

rebeccaxhawaii (rebeccaxhawaii):

D is correct

OpenStudy (anonymous):

The regular price of an item is $65. The item is on sale with a discount of $8. 12.3% is the discount rate. 100 x 8/65 = 12.3